# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v23

@@ 3.18.1 Printed Text
The XTREME-UP120120120https://github.com/google-research/xtreme-up data set created by Ruder et al. [137] contains a Sinhala data set for the OCR task. The data was obtained from book transcriptions. An attempt for Sinhala OCR system has been taken by Rajapakse et al. [387] before any other work has been done on the topic. Much later, a linear symmetry-based approach was proposed by Premaratne and Bigun [388, 389]. They then used hidden Markov model-based optimization on the recognized Sinhala script [390]. Similarly, Hewavitharana et al. [391] used hidden Markov models for off-line Sinhala character recognition. Herath et al. [119], Herath and Medagoda [120] developed a prepossessing engine based on a template matching for printed Sinhala OCR. Statistical approaches with histogram projections for Sinhala character recognition is proposed by Hewavitharana and Kodikara [392], by Ajward et al. [393], and by Madushanka et al. [394].  Karunanayaka et al. [395] also did off-line Sinhala character recognition with a use case for postal city name recognition. A separate group had attempted Sinhala OCR [396] mainly involving the nearest-neighbor method [397, 398]. A study by Ediriweera [399] uses dictionaries to correct errors in Sinhala OCR. An early attempt for Sinhala OCR by Dias et al. [400] has been extended to be online and made available to use via desktops [401] and hand-held devices [402] with the ability to recognize handwriting. A simple neural network based approach for Sinhala OCR was utilized by Rimas et al. [403]. A fuzzy-based model for identifying printed Sinhala characters was proposed by Gunarathna et al. [404]Premachandra et al. [405] proposes a simple back-propagation artificial neural network with hand crafted features for Sinhala character recognition. Another neural network with specialized feature extraction for Sinhala character recognition was proposed by Jayamaha and Naleer [406]. On the matter of neural networks and feature extraction, a feature selection process for Sinhala OCR was proposed by Kumara and Ragel [407].  Jayawickrama et al. [408] worked on Sinhala printed characters with special focus on handling diacritic vowels. However, they opted to refer to diacritic vowels as modifiers in their work. Gunawardhana and Ranathunga [409] proposed a limited approach to recognize Sinhala letters on Facebook images. A CNN-based methodology to improve printed Sinhala character OCR was proposed by Liyanage [410]. Printed Character Recognition (PCR) was used by Vasantharajan and Thayasivam [109] to create a large-scale Tamil-Sinhala-English parallel corpus. A meta-study on the effects of text genre, image resolution, and algorithmic complexity needed for Sinhala OCR from books and newspapers was conducted by Anuradha et al. [411].  Anuradha et al. [412] used Tesseract 3121121121https://tesseract-ocr.github.io/ [413] for Sinhala OCR. The later study by [414] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.  Maduranga and Jayalal [415] used Artificial Neural Network (ANN) based on Universe of Discourse and Self Organization Map methods to recognize multi-style printed Sinhala characters. The study by De Silva et al. [416], even though ostensibly presented as translation research, is just invoking Google Cloud service for translation. Their novelty in Sinhala NLP lies in the facility provided to OCR the text from photos and documents. The work by Thamarasee and Surendra [417] used histogram-oriented gradient descriptor (HOG descriptor) [418] and SVMs to recognize 400+ variations of Sinhala characters.
@@ 3.18.3 Ancient Text
Summarizing optically recognized old Sinhala text for the purpose of archival search and preservation was explored by Rathnasena et al. [443]. The work of Peiris [444] also focused on OCR for ancient Sinhala inscriptions. Building upon the architecture proposed by Ruwanmini et al. [445], a neural network-based method for recognizing ancient Sinhala inscriptions was proposed by Karunarathne et al. [446]. The study by Wickramarathna and Ranathunga [447] created a system to recognize Brahmi characters, correct errors, and generate Sinhala meanings. Heenkenda and Fernando [448] used Inception-v3 [449], VGG-19 [450], and ResNet-50 [451] to classify Sinhala inscriptions to historical time periods. The work by Gunasekara et al. [452] claims to have created a mobile app that is capable of recognizing and translating early Brahmi characters.
@@ 3.19.2 Sinhala-English NMT
Fonseka et al. [476] used Byte Pair Encoding (BPE) for English to Sinhala neural machine translation. As another solution to the OOV problem, an analysis of subword techniques to improve English to Sinhala Neural Machine Translation (NMT) was conducted by Naranpanawa et al. [477]. A data augmentation method to expand bilingual lexicon terms based on case markers for the purpose of solving the OOV problem in the domain of NMT was proposed by Fernando et al. [478] which they later extended further [479]Epaliyana et al. [480] proposed iterative filtering and data selection be used to improve Sinhala-English NMT. Perera et al. [481] used English Part-of-Speech (PoS) tags to improve English to Sinhala NMT. Lin et al. [482] used a model based on fairseq [483] to improve machine translation between English and Sinhala. The second half of the study by Arafath [360] dealt with translating Sinhala speech to other languages. Kugathasan and Sumathipala [484] proposed an NMT system for Sinhala-English Code-Mixed text using the standardized Sinhala Code-Mixed text they proposed earlier [485]. Nguyen et al. [486] introduced a new fine-tuning objective LAgSwAV (Language-Agnostic Constraint for SwAV loss), using which they obtained 5.4 BLEU for English-Sinhala. Later, this method was further discussed by Nguyen [487]. Attigala and Weerasinghe [488] conducted an analysis on the effectiveness of ChatGPT [489] in translating Sinhala songs to English. Utsa et al. [490] used transfer learning and back translation as well as focal loss of the Sinhala-English dataset from FLoRes [95] and reported better results than mBART [491].
@@ 3.19.3 Singlish to Sinhala (Transliteration)
The XTREME-UP122122122https://github.com/google-research/xtreme-up data set created by Ruder et al. [137] contains a Sinhala data set for the transliteration task. The early work by Goonetilleke et al. [492] attempted Sinhala transliteration through the Latin alphabet. However, their work does not use the word transliteration and instead focuses on the predictive aspect. Priyadarshani et al. [493] used statistical machine learning for transliteration of names between Sinhala, Tamil, and English. A rule-based system on trigrams was proposed by Liwera and Ranathunga [494] for Singlish to Sinhala transliteration of social media text. A Singlish to Sinhala converter which uses an LSTM was proposed by De Silva [495]. A rule-based approach for the same was proposed by de Silva and Ahangama [496]. The study by Nanayakkara et al. [497] introduced an English-to-Sinhala transliteration system. Sumanathilaka et al. [498, 499] proposed a Trie [500] data structure-based algorithm for word suggestion in Sinhala transliterations named Swa-Bhasha. An extended analysis of the same work was presented in a later work [501]. A further extension, Swa-Bhasha 2.0, by Dharmasiri and Sumanathilaka [502] used NMT to eliminate word selection ambiguity in the transliteration process. Amarasekara et al. [503] proposed a rule-based method supported by N-gram analysis and a corpus dataset to transliterate Singlish tweets to Sinhala. The study by Rajapaksha et al. [504] claims to have created a translation system, however, what they have created is a transliteration system between Singlish and Sinhala. While they also claim to have trained ASR and TTS systems, the paper does not identify them by name or citation. The same is true for the data sets they have used to train the various modules in their pipeline. The studies by Kumaravithana et al. [505] and Jayawardhana et al. [506] claim to have built Sinhala-English translators but in actuality are just invoking Google API for their translation tasks.
@@
The work by Kirov et al. [507] used the Dakshina [150] dataset and compared the accuracy of Singlish to Sinhala transliteration with vanilla LSTM, vanilla Transformers, mT5 [310], ByT5 [508], and a non-neural finite-state transducer (FST) based on work by Bisani and Ney [509]. They discuss in detail how the zero-width joiner (ZWJ, U+200D) character cause issues in this task and Sinhala rendering as a whole. Their code is available on GitHub123123123https://github.com/google-research/google-research/tree/master/context_aware_transliteration. Khiu et al. [510] fine-tuned mBART on Sinhala Government Corpus and Bible corpus to predict the NMT performance dependence on domain similarity.
@@ 3.19.5 Between Sinhala and Non-English Languages
Most of the cross-Sinhala and Tamil work has been done in the domain of machine translation. A neural machine translation for Sinhala and Tamil languages was initiated by Tennage et al. [514, 515]. Then they further enhanced it with transliteration and byte pair encoding [516] and used synthetic training data to handle the rare word problem [517]. This project produced Si-Ta [518] a machine translation system of Sinhala and Tamil official documents. In the statistical machine translation front, Farhath et al. [519] worked on integrating bilingual lists. The attempts by Weerasinghe [520] and Sripirakas et al. [521] were also focused on statistical machine translation while Jeyakaran and Weerasinghe [522] attempted a kernel regression method. A yet another attempt was made by Pushpananda et al. [523] which they later extended with some quality improvements [524]. An attempt at real-time direct translation between Sinhala and Tamil was done by Rajpirathap et al. [525]Dilshani et al. [526] have done a study on the linguistic divergence of Sinhala and Tamil languages with respect to machine translation. Mokanarangan [527] claims to have built a named entity translator between Sinhala and Tamil for official government documents. But this work is locked behind an institutional repository wall and thus is not accessible by other researchers. Arukgoda et al. [528] studied the possibility of using deep learning techniques to improve Sinhala-Tamil translation which they further improved later [529]Pramodya et al. [530] compared Transformers, Recurrent Neural Networks, and Statistical Machine Translation (SMT) in the context of Tamil to Sinhala machine translation. The work by Nissanka et al. [531] used monolingual word embedding to improve NMT between Sinhala and Tamil. Thillainathan et al. [532] uses pre-trained mBART [491] models for six directional translations between Sinhala, Tamil, and English. Yashothara and Uthayasanker [533] discussed the use of the Hierarchical Phrase-Based Model for Tamil to Sinhala and Sinhala to Tamil translations. Pramodya [534] presented a comparison of SMT [473] and NMT models for Sinhala-Tamil translation. For NMT they have used base transformer, turned transformer, and mT5 [310]. The work by Su et al. [535] which compared 8 Parameter-efficient fine-tuning (PEFT) methods on the Si-Ta language pair using the Government corpus [474] and NLLB [130] reports that Houlsby adapter [233], with a 33.34 SacreBLEU [536] score, to be the best PEFT method. However, they also note that the Pfeiffer adapter [235] runs the fastest at 52.59 hours with a reasonable score of 31.24. (Comparatively, the Houlsby adapter runs for 78.65 hours). The work by Pramodya et al. [537] conducted a comparative analysis of the available models for Sinhala-Tamil translation and reaffirmed the observations of Duh et al. [538] that, in low-resource scenarios, SMT and NMT both may work similarly, however, NMT needs more careful tuning to fit performance.
@@
There have been attempts to link Sinhala NLP with Japanese by Herath et al. [23, 90, 89]Thelijjagoda et al. [91]Thelijjagoda et al. [539], and Kanduboda [10]. Jayasinghe [540] discusses the importance of translation between Sinhala and the liturgical language of Buddhism, Pali [541, 542, 543]. [544] attempted to use a dictionary-based machine-translation method for this task. The study by Anuradha and Thelijjagoda [336] uses machine translation on the unique application of converting Sinhala and English Braille documents, which they have run OCR on, into voice.
@@ 3.20 Spelling and Grammar
The open-source data-driven approach proposed by Wasala et al. [545, 546] claims to be able to check and correct spelling errors in Sinhala. The approach by Jayalatharachchi et al. [547] attempts to obtain synergy between two algorithms for the same purpose. These efforts [545, 547] were then extended by Subhagya et al. [548]. A rule-based Sinhala spell checker named SinSpell based on Hunspell124124124http://hunspell.github.io/ was introduced by Liyanapathirana et al. [549]. They have also made the tool available125125125http://nlp-tools.uom.lk/sinspell/ online for use. The study by Sithamparanathan and Uthayasanker [550] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil. Sonnadara et al. [123] created a benchmark data set for Sinhala spell correction along with a neural model. A multi agent-based spell checker, named LaSi Spell for Sinhala spell checking was introduced by Samarawickrama et al. [551]. The study by Udagedara et al. [552] specifically solved the problem of spell-checking Sri Lankan names and addresses. A system named Erroff was proposed by Sudesh et al. [553] to correct real-word errors in Sinhala text.

### v24

@@ 3.17.6 Lip Synchronization
3.17.7 Music to Notation
The work by Dulmi and Hettige [417] has created a pipeline which converts audio recordings to western music notation which is then converted to Sinhala musical notations using an API call to a Large Language Model with a custom-tailored prompt.
@@ 3.18.1 Printed Text
The XTREME-UP140140140https://github.com/google-research/xtreme-up data set created by Ruder et al. [144] contains a Sinhala data set for the OCR task. The data was obtained from book transcriptions. An attempt for Sinhala OCR system has been taken by Rajapakse et al. [418] before any other work has been done on the topic. Much later, a linear symmetry-based approach was proposed by Premaratne and Bigun [419, 420]. They then used hidden Markov model-based optimization on the recognized Sinhala script [421]. Similarly, Hewavitharana et al. [422] used hidden Markov models for off-line Sinhala character recognition. Herath et al. [123], Herath and Medagoda [124] developed a prepossessing engine based on a template matching for printed Sinhala OCR. Statistical approaches with histogram projections for Sinhala character recognition is proposed by Hewavitharana and Kodikara [423], by Ajward et al. [424], and by Madushanka et al. [425].  Karunanayaka et al. [426] also did off-line Sinhala character recognition with a use case for postal city name recognition. A separate group had attempted Sinhala OCR [427] mainly involving the nearest-neighbor method [428, 429]. A study by Ediriweera [430] uses dictionaries to correct errors in Sinhala OCR. An early attempt for Sinhala OCR by Dias et al. [431] has been extended to be online and made available to use via desktops [432] and hand-held devices [433] with the ability to recognize handwriting. A simple neural network based approach for Sinhala OCR was utilized by Rimas et al. [434]. A fuzzy-based model for identifying printed Sinhala characters was proposed by Gunarathna et al. [435]Premachandra et al. [436] proposes a simple back-propagation artificial neural network with hand crafted features for Sinhala character recognition. Another neural network with specialized feature extraction for Sinhala character recognition was proposed by Jayamaha and Naleer [437]. On the matter of neural networks and feature extraction, a feature selection process for Sinhala OCR was proposed by Kumara and Ragel [438].  Jayawickrama et al. [439] worked on Sinhala printed characters with special focus on handling diacritic vowels. However, they opted to refer to diacritic vowels as modifiers in their work. Gunawardhana and Ranathunga [440] proposed a limited approach to recognize Sinhala letters on Facebook images. A CNN-based methodology to improve printed Sinhala character OCR was proposed by Liyanage [441]. Printed Character Recognition (PCR) was used by Vasantharajan and Thayasivam [113] to create a large-scale Tamil-Sinhala-English parallel corpus. A meta-study on the effects of text genre, image resolution, and algorithmic complexity needed for Sinhala OCR from books and newspapers was conducted by Anuradha et al. [442].  Anuradha et al. [443] used Tesseract 3141141141https://tesseract-ocr.github.io/ [444] for Sinhala OCR. The later study by [445] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.  Maduranga and Jayalal [446] used Artificial Neural Network (ANN) based on Universe of Discourse and Self Organization Map methods to recognize multi-style printed Sinhala characters. The study by De Silva et al. [447], even though ostensibly presented as translation research, is just invoking Google Cloud service for translation. Their novelty in Sinhala NLP lies in the facility provided to OCR the text from photos and documents. The work by Thamarasee and Surendra [448] used histogram-oriented gradient descriptor (HOG descriptor) [449] and SVMs to recognize 400+ variations of Sinhala characters.
@@ 3.18.3 Ancient Text
Summarizing optically recognized old Sinhala text for the purpose of archival search and preservation was explored by Rathnasena et al. [474]. The work of Peiris [475] also focused on OCR for ancient Sinhala inscriptions. Building upon the architecture proposed by Ruwanmini et al. [476], a neural network-based method for recognizing ancient Sinhala inscriptions was proposed by Karunarathne et al. [477]. The study by Wickramarathna and Ranathunga [478] created a system to recognize Brahmi characters, correct errors, and generate Sinhala meanings. Heenkenda and Fernando [479] used Inception-v3 [480], VGG-19 [481], and ResNet-50 [482] to classify Sinhala inscriptions to historical time periods. The work by Gunasekara et al. [483] claims to have created a mobile app that is capable of recognizing and translating early Brahmi characters. The work by Surasinghe and Thanikasalam [484] used CNN models to predict the historical period of ancient Sinhala text. In their subsequent work [485], they extended this work by adding attention to the CNN models.
@@ 3.19.2 Sinhala-English NMT
Fonseka et al. [509] used Byte Pair Encoding (BPE) for English to Sinhala neural machine translation. As another solution to the OOV problem, an analysis of subword techniques to improve English to Sinhala Neural Machine Translation (NMT) was conducted by Naranpanawa et al. [510]. A data augmentation method to expand bilingual lexicon terms based on case markers for the purpose of solving the OOV problem in the domain of NMT was proposed by Fernando et al. [511] which they later extended further [512]Epaliyana et al. [513] proposed iterative filtering and data selection be used to improve Sinhala-English NMT. Perera et al. [514] used English Part-of-Speech (PoS) tags to improve English to Sinhala NMT. Lin et al. [515] used a model based on fairseq [516] to improve machine translation between English and Sinhala. The second half of the study by Arafath [387] dealt with translating Sinhala speech to other languages. Kugathasan and Sumathipala [517] proposed an NMT system for Sinhala-English Code-Mixed text using the standardized Sinhala Code-Mixed text they proposed earlier [518]. Nguyen et al. [519] introduced a new fine-tuning objective LAgSwAV (Language-Agnostic Constraint for SwAV loss), using which they obtained 5.4 BLEU for English-Sinhala. Later, this method was further discussed by Nguyen [520]. Attigala and Weerasinghe [521] conducted an analysis on the effectiveness of ChatGPT [522] in translating Sinhala songs to English. Utsa et al. [523] used transfer learning and back translation as well as focal loss of the Sinhala-English dataset from FLoRes [98] and reported better results than mBART [524]. The study by Ranathungaa et al. [525] presents an evaluation of using parallel data from auxiliary domains to enhance EnSI NMT by fine-tuning or further pre-training the models. In their study on improving cross-lingual representation of multilingual language models for low-resource languages using linguistic entity masking, Fernando and Ranathunga [288] used SiTa and EnSi NMT on NLLB and CCAligned as one of the evaluations.
Sinhala is one of the languages considered in the short survey by Perera and Sumanathilaka [526] on machine translation and transliteration for Indo-Aryan Languages. They mention a sub-set of works [505, 509, 527, 508, 528, 529] discussed in this survey for EnSi translation.
@@ 3.19.3 Singlish to Sinhala (Transliteration)
The XTREME-UP142142142https://github.com/google-research/xtreme-up data set created by Ruder et al. [144] contains a Sinhala data set for the transliteration task. The early work by Goonetilleke et al. [530] attempted Sinhala transliteration through the Latin alphabet. However, their work does not use the word transliteration and instead focuses on the predictive aspect. Priyadarshani et al. [531] used statistical machine learning for transliteration of names between Sinhala, Tamil, and English. A rule-based system on trigrams was proposed by Liwera and Ranathunga [532] for Singlish to Sinhala transliteration of social media text. A Singlish to Sinhala converter which uses an LSTM was proposed by De Silva [533]. A rule-based approach for the same was proposed by de Silva and Ahangama [534]. The study by Nanayakkara et al. [535] introduced an English-to-Sinhala transliteration system.
Athukorala and Sumanathilaka [536] proposed a fuzzy logic-based Sinhala transliteration system named Swa-Bhasha. Sumanathilaka et al. [537, 538] then proposed a Trie [539] data structure-based algorithm for word suggestion for the Swa-Bhasha system. An extended analysis of the same work was presented in a later work [540]. A further extension, Swa-Bhasha 2.0, by Dharmasiri and Sumanathilaka [541] used NMT to eliminate word selection ambiguity in the transliteration process. Subsequently, Sumanathilaka et al. [542] claim to have released the Swa-Bhasha Dataset.
Amarasekara et al. [543] proposed a rule-based method supported by N-gram analysis and a corpus dataset to transliterate Singlish tweets to Sinhala. The study by Rajapaksha et al. [544] claims to have created a translation system, however, what they have created is a transliteration system between Singlish and Sinhala. While they also claim to have trained ASR and TTS systems, the paper does not identify them by name or citation. The same is true for the data sets they have used to train the various modules in their pipeline. The studies by Kumaravithana et al. [545] and Jayawardhana et al. [546] claim to have built Sinhala-English translators but in actuality are just invoking Google API for their translation tasks.
@@
The work by Kirov et al. [547] used the Dakshina [157] dataset and compared the accuracy of Singlish to Sinhala transliteration with vanilla LSTM, vanilla Transformers, mT5 [336], ByT5 [548], and a non-neural finite-state transducer (FST) based on work by Bisani and Ney [549]. They discuss in detail how the zero-width joiner (ZWJ, U+200D) character cause issues in this task and Sinhala rendering as a whole. Their code is available on GitHub143143143https://github.com/google-research/google-research/tree/master/context_aware_transliteration. Khiu et al. [550] fine-tuned mBART on Sinhala Government Corpus and Bible corpus to predict the NMT performance dependence on domain similarity. De Mel et al. [551] conducted a comparative analysis for Sinhala transliteration between an extension of the rule-based approach proposed by Tennage et al. [552] and the transliteration as a translation model proposed by Deselaers et al. [553]. They reported across-the-board better performance with the latter. All thier code as well as the test-sets are publicly avaialble144144144https://github.com/kasunw22/Sinhala-Transliterator/. Perera et al. [554] attempted the same task with BERT. Their code is also publicly available145145145https://github.com/Sameera2001Perera/Singlish-Transliterator.
In their short survey discussing Indo-Aryan Languages machine translation and transliteration, Perera and Sumanathilaka [526] covers a number of works [504, 552, 531, 532, 534, 536, 535, 538, 541, 542] on Sinhala transliteration which we have already discussed above.
@@ 3.19.5 Between Sinhala and Non-English Languages
Most of the cross-Sinhala and Tamil work has been done in the domain of machine translation. A neural machine translation for Sinhala and Tamil languages was initiated by Tennage et al. [557, 558]. Then they further enhanced it with transliteration and byte pair encoding [552] and used synthetic training data to handle the rare word problem [559]. This project produced Si-Ta [527] a machine translation system of Sinhala and Tamil official documents. In the statistical machine translation front, Farhath et al. [560] worked on integrating bilingual lists. The attempts by Weerasinghe [561] and Sripirakas et al. [562] were also focused on statistical machine translation while Jeyakaran and Weerasinghe [563] attempted a kernel regression method. A yet another attempt was made by Pushpananda et al. [564] which they later extended with some quality improvements [565]. An attempt at real-time direct translation between Sinhala and Tamil was done by Rajpirathap et al. [566]Dilshani et al. [567] have done a study on the linguistic divergence of Sinhala and Tamil languages with respect to machine translation. Mokanarangan [568] claims to have built a named entity translator between Sinhala and Tamil for official government documents. But this work is locked behind an institutional repository wall and thus is not accessible by other researchers. Arukgoda et al. [569] studied the possibility of using deep learning techniques to improve Sinhala-Tamil translation which they further improved later [570]Pramodya et al. [571] compared Transformers, Recurrent Neural Networks, and Statistical Machine Translation (SMT) in the context of Tamil to Sinhala machine translation. The work by Nissanka et al. [572] used monolingual word embedding to improve NMT between Sinhala and Tamil. Thillainathan et al. [573] uses pre-trained mBART [524] models for six directional translations between Sinhala, Tamil, and English. Yashothara and Uthayasanker [574] discussed the use of the Hierarchical Phrase-Based Model for Tamil to Sinhala and Sinhala to Tamil translations. Pramodya [575] presented a comparison of SMT [506] and NMT models for Sinhala-Tamil translation. For NMT they have used base transformer, turned transformer, and mT5 [336]. The work by Su et al. [576] which compared 8 Parameter-efficient fine-tuning (PEFT) methods on the Si-Ta language pair using the Government corpus [507] and NLLB [134] reports that Houlsby adapter [254], with a 33.34 SacreBLEU [577] score, to be the best PEFT method. However, they also note that the Pfeiffer adapter [256] runs the fastest at 52.59 hours with a reasonable score of 31.24. (Comparatively, the Houlsby adapter runs for 78.65 hours). The work by Pramodya et al. [578] conducted a comparative analysis of the available models for Sinhala-Tamil translation and reaffirmed the observations of Duh et al. [579] that, in low-resource scenarios, SMT and NMT both may work similarly, however, NMT needs more careful tuning to fit performance. The short survey [526] on Indo-Aryan Languages machine translation and transliteration, discusses a few works [527, 572, 509, 571, 569, 573] under SiTa translation which have already been discussed in this survey.
@@
There have been attempts to link Sinhala NLP with Japanese by Herath et al. [27, 95, 94]Thelijjagoda et al. [96]Thelijjagoda et al. [580], and Kanduboda [10]. Jayasinghe [581] discusses the importance of translation between Sinhala and the liturgical language of Buddhism, Pali [582, 583, 584]. Shalini and Hettige [585] attempted to use a dictionary-based machine-translation method for this task. The study by Anuradha and Thelijjagoda [363] uses machine translation on the unique application of converting Sinhala and English Braille documents, which they have run OCR on, into voice.
@@ 3.20 Spelling and Grammar
The open-source data-driven approach proposed by Wasala et al. [586, 587] claims to be able to check and correct spelling errors in Sinhala. The approach by Jayalatharachchi et al. [588] attempts to obtain synergy between two algorithms for the same purpose. These efforts [586, 588] were then extended by Subhagya et al. [589]. A rule-based Sinhala spell checker named SinSpell based on Hunspell146146146http://hunspell.github.io/ was introduced by Liyanapathirana et al. [590]. They have also made the tool available147147147http://nlp-tools.uom.lk/sinspell/ online for use. The study by Sithamparanathan and Uthayasanker [591] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil. Sonnadara et al. [127] created a benchmark data set for Sinhala spell correction along with a neural model. A multi agent-based spell checker, named LaSi Spell for Sinhala spell checking was introduced by Samarawickrama et al. [592]. The study by Udagedara et al. [593] specifically solved the problem of spell-checking Sri Lankan names and addresses. A system named Erroff was proposed by Sudesh et al. [594] to correct real-word errors in Sinhala text. Praveen et al. [362] claim to have developed a Sinhala Spell checker with the Random Forest algorithm that archives 82% accuracy.